# Improving Code Reviewer Recommendation: Accuracy, Latency, Workload, and Bystanders

### v1

@@ Abstract.
Code review ensures that a peer engineer manually examines the code before it is integrated and released into production. At Meta, we develop a wide range of software at scale, from social networking to software development infrastructure, such as calendar and meeting tools to continuous integration. We are constantly improving our code review system, and in this work we describe a series of experiments that were conducted across 10’s of thousands of engineers and 100’s of thousands of reviews.
We build upon the recommender that has been in production since 2018, RevRecV1. We found that reviewers were being assigned based on prior authorship of files. We reviewed the literature for successful features and experimented with them with RevRecV2 in production. The most important feature in our new model was the familiarity of the author and reviewer, we saw an overall improvement in accuracy of 14 percentage points.
Prior research has shown that reviewer workload is skewed. To balance workload, we divide the reviewer score from RevRecV2 by each candidate reviewers workload. We experimented with multiple types of workload to develop RevRecWL. We find that reranking candidate reviewers by workload often leads to a reviewers with lower workload being selected by authors.
The bystander effect can occur when a team of reviewers is assigned the review. We mitigate the bystander effect by randomly assigning one of the recommended reviewers. Having an individual who is responsible for the review, reduces the time take for reviews by -11%.
@@ 1. Introduction
Code review is critical activity in the software development lifecycle that ensures that at least one other experienced developer examines the code for defects and other issues. Code review has evolved from the inspection of massive completed artifacts (Fagan, 1976) to the continuous code review of the difference between the changed files in the form of pull-requests, patchsets, and diffs (Rigby and Bird, 2013; Bacchelli and Bird, 2013).
Expertise has been repeatedly shown to be a key factor in good code reviews (Porter et al., 1998; Rigby et al., 2014). Reviewer recommenders have been developed to suggest candidate reviewers who are most likely to perform high quality reviews, e.g., (Balachandran, 2013; Thongtanunam et al., 2015). Much of this work has been based on experience, e.g., this person has reviewed and authored the files in a change in the past. Other types of features including the relationships among developers have been used to identify candidate reviewers, e.g., (Jiang et al., 2017; Zhang et al., 2022). Most works have only tried to solve the problem of finding an expert reviewer, however, recent works have started to consider other factors such as knowledge spreading (Mirsaeedi and Rigby, 2020) and reviewer workload (Asthana et al., 2019; Strand et al., 2020; Al-Zubaidi et al., 2020)
@@
At Meta, we are constantly evolving our code review system. In this work, we describe a series of improvements that we made to our code reviewer recommenders. Our starting point was developed in 2018 and we refer to it as RevRecV1. To address latency and accuracy issues with RevRecV1, we developed RevRecV2 that is based on a literature review of successfully used features. Building on RevRecV2, we use measures of reviewer workload to spread review work and develop RevRecWL. Finally, we combat the bystander effect when a group is assigned to do a review with BystanderRecRnd. One of our major contributions is to experiment with our recommenders in production and to conduct A vs B double blind trials on each recommender across hundreds of thousands of reviews and 10’s of thousands of engineers. We provide an overview of the motivation for each recommender and briefly summarize the results.
@@
1. Accuracy and latency improvements with RevRecV2. On the code review team’s feedback group, some engineers complained that they were being incorrectly recommended for reviews. When we investigated, we found that the current recommender, RevRecV1, was purely based on hg blame lines and would suggest reviewers who had moved onto other projects because they had edited the files in the past. We reviewed the literature and developed a recommender that took the most promising features into account.
@@
Result summary: In an A/B test, we saw a top 3 accuracy of 73% a 14 point improvement over the previous model. We also saw a 14 times reduction in latency at the 90th percentile. The recommendations were selected by authors 27% more often with the new model. We did not see any statistically significant changes in the review cycle time or the reviewer viewing time. In Summer 2022, RevRecV2 was rolled out to 100% of engineers.
@@
2. Reviewer workload balancing with RevRecWL. The number of reviews that each engineer does is known to be skewed (Rigby et al., 2014; Hajari, 2022; Al-Zubaidi et al., 2020). At Microsoft (Asthana et al., 2019) and Ericsson (Strand et al., 2020), they implemented a recommender that weighted the expertise based recommender score by the workload of the candidate reviewer. We replicate these works and experiment with novel workload measures, including the number of scheduled meetings to understand if workload balancing is effective.
@@
Result summary: When a reasonable candidate with lower workload was available, we saw a large reduction in workload. When an inappropriate candidate was recommended, simply because they had low workload, the author looked down the list for a more experienced developer, and we saw a drop in the top ranked candidate being selected. We did not see any statistically significant change in review cycle time or latency. In Winter of 2023, RevRecWL was rolled out to 100% of engineers.
@@
3. Reducing the bystander effect with BystanderRecRnd. The bystander effect occurs when nobody is explicitly assigned responsibility for a situation and no individual takes action or they are delayed in taking action (Fischer et al., 2011). Beyond noting that it may be an issue in code review (Meneely et al., 2014), we did not find any research in the software engineering literature. At Meta teams and groups of developers can be assigned a review without any explicit individual. To reduce the bystander effect, we use our recommenders to explicitly assign a reviewer.
@@
Result summary: At Meta, we randomly assigned one of the top3 experts to review the diff and found a drop of -11.6% in time-in-review with no regressions in our guardrail metrics. In Spring of 2023, we rolled BystanderRecRnd out to 100% of engineers.
This works is structured contributions:
  1. (1)
    In Section 2, we describe Meta and its review process.
  2. (2)
    In Section 3, we review the literature to identify the features that have been successfully used to recommend reviewers.
  3. (3)
    In Section 5, we update the the recommender at Meta to include these features and experiment with a new model in production.
  4. (4)
    In Section 6, we weight the recommender by the workload of the candidate reviewers in an attempt to balance workload.
  5. (5)
    In Section 7, we reduce the impact of bystander effect when no individual is assigned a review by randomly assigning one of the recommended reviewers
  6. (6)
    We conduct A/B tests on each recommender including 100’s of thousands of reviews and 10’s of thousands of engineers across a diverse range of software from web infrastructure and social networking to internal tooling including continuous integration systems.
In the final sections of the paper, we discuss threats to validity in Section 8, position our work in the literature in Section 9, and conclude the paper in Section 10.
@@ 2. Background and Code Review Process
Meta is a large online services company that works across the spectrum in the communications, retail, entertainment industries. Meta has tens of thousands of employees with offices spread across the globe (North America, Europe, Middle East and Asia). Meta has its own dedicated infrastructure teams where the tools used are a mix of commercial and in-house developed systems. We study code reviews from the projects using Phabricator  [^1] Phabricator is an open source project available at https://www.phacility.com/phabricator/ which includes user facing social network products and virtual and augmented reality projects as well as software engineering infrastructure, such as calendar, tasks, and release engineering tooling.
@@ 2.1. Code review process
Code review at Meta follows a similar structure to the contemporary practices used at Google (Sadowski et al., 2018) and Microsoft (Bacchelli and Bird, 2013), and we briefly summarize the practice here. Figure 1 provides an example of a diff review with portions redacted and using mock names. Like most major code review tools, the code author provides a title and a detailed summary of the change. A test plan is provided that details which steps the author took to validate their changes, and provides reviewers a guide to test the change themselves. Line-by-line highlighting is used to show additions and removals, and inline comments are used to indicate any compilation or linting issues. The diff will also show the output of tests that were required to run based on the files included in the change.
Before publishing a code change, authors select one or more reviewers and/or reviewer groups. Reviewers provide feedback, ask questions and make suggestions, which should be addressed by the author. There may be multiple rounds of feedback and changes before a reviewer accepts the diff. Meta uses a single track/monorepo development model, so once the diff is accepted it will be integrated for further testing and ultimately “shipped“ into production.
@@ 2.2. Reviewer Recommenders at Meta
The reviewer recommender used in production at Meta before we began our work was developed in 2018. For this study it will be our starting comparison point and we will refer to it as RevRecV1. RevRecV1 considered the following attributes: the blame lines i.e. the person who has authored the most lines in a file is an expert, authorship frequency i.e. the person who has made the most changes to a file is an expert, and review frequency i.e. the number of times a reviewer has accepted the changes to a file. RevRecV1 is trained using a genetic algorithm, an ML approach popular in 2018. While the recommender has been effective and makes thousands of recommendations every day, the code review team identified a number of issues with it. First, many developers complain about incorrect recommendations on the feedback group. Second, hg-blame is an computationally expensive command that can delay the loading of the code review page for a diff. For these reasons, the code review team decided to test alternative strategies for recommending reviewers to improve accuracy and reduce latency.
3. Literature and Features
We reviewed the literature to identify the reviewer recommendation approaches and model features that were most successful in existing recommenders. Like Microsoft (Asthana et al., 2019; Zanjani et al., 2016) and Ericsson (Strand et al., 2020), we cannot use features in production that would add latency to our recommender. We also found that very few recommenders that are studied in the literature report latency and even fewer are actually deployed into production. We break the literature into the following categories (Al-Zubaidi et al., 2020): code ownership, reviewing experience, review participation rate, and familiarity with the author. Table 2 shows the features we selected for RevRecV2.
@@ 3. Literature and Features
The influence of code ownership on code quality has been extensively investigated in the literature(Bird et al., 2011; Rahman and Devanbu, 2011; Foucault et al., 2015; Thongtanunam et al., 2016) with recommenders to identify expert developers for a particular part of code or task (Mockus and Herbsleb, 2002). Researchers have used a wide range of granularity, from lines (Fritz et al., 2007; Girba et al., 2005; Rahman and Devanbu, 2011) to modules (Bird et al., 2011), to estimate ownership of developers. Studies on code review find that code owners are usually selected to review changes (Sadowski et al., 2018; Bacchelli and Bird, 2013; Greiler et al., 2016). In this work, we quantify ownership as the number of files under review that a candidate reviewer has modified in the past diffs.
@@
Thongtanunam et al. (Thongtanunam et al., 2016) recommend reviewers based on the files that an engineer has reviewed. Intuitively, engineers who have reviewed the files in the diff in the past, will be good candidate reviewers. Zanjani et al. (Zanjani et al., 2016) build upon this notion of reviewer expertise by adding the number of comments a candidate reviewer has made on the files under review. In this work, we count the number of comments as well as other reviewer actions such as accepting or rejecting the diff.
@@
Moving beyond code ownership, reviewer recommenders have begun to consider relationships among the reviewer and the author (Bosu et al., 2016). The simplest metric is the number of times a reviewer and has commented on an authors diffs in the past. More sophisticated usages include social relationship graphs (Yu et al., 2016). With 10’s of thousands of developers, and thousands of diffs every day, social network graphs are not performant. In this work, we suggest a novel usage of how long a reviewer looks at an authors diffs: the time a reviewer spent looking at the diffs the author wrote in the code review tool. While Meta has sophisticated tracking of how long an engineer is looking at a webpage and uses this to capture how long a reviewer looks at a review, other projects can easily add this to the review tool. For example, the SmartBear code review tool has been tracking this information since 2006 (Cohen, 2006) to identify reviewers that looked at the change for too short a period of time given the size of the diff. The web abounds with code to monitor time spent on a webpage. [^2] A JavaScript library that tracks time spent on a page https://stackoverflow.com/questions/4667068/how-to-measure-a-time-spent-on-a-page
@@
Recent works have incorporated the reviewer workload into reviewer recommendation. Early works quantified the relationship between time waiting for a review and review workload. Baysal et al. (Baysal et al., 2013) studied WebKit to understand the impact of metrics on review time. They found that when reviewers had longer queues of waiting reviews, reviews tended to take longer. They also found that more active reviewers had quicker response times than less active reviewers. Studying four open source projects, Ruangwan et al. (Ruangwan et al., 2019) found prior participation and prior experience were the best predictor of whether a reviewer will respond to a review request. The familiarity of the author an reviewer had a minor impact on participation.
@@
Microsoft weighted their code reviewer recommender, cHRev, by the number of open reviews a candidate, WhoDo, had and found between a 14% and 21% improvement in the time taken for reviews depending on the project (Asthana et al., 2019). In contrast, Strand  et al. (Strand et al., 2020) evaluate a workload aware recommender at Ericsson. In a backtest they found that a recommender aware of the file change history and the level of recent workload activity was able to make recommendations with an MRR of .37 or an average rank of 2.7. Despite these reasonable recommendations, after manually evaluating the performance of the recommender on 47 changes they determined that it did not reduce lead time or workload for code review and they decided to not rollout the recommender.
4. Recommender Design and Evaluation Metrics
Table 1. We report accuracy, and we also have goal and guardrail metrics which we discuss for each experiment.
Outcome Metric Description
TopN Accuracy
How often does one of the recommended reviewers review the diff? We display 5 reviewers and report Top1 and Top3 accuracy.
TimeInReview
The time when a diff is waiting for review or is being actively reviewed. It excludes the time when the author is reworking the code. We want to make sure that our recommender does not slow down reviewers.
TimeSpent
The amount of time reviewers spend examining each diff. This is time spent on the review page for a diff. We want to make sure that reviewers are not rushing through reviews and rubberstamping diffs.
Latency
The amount of time from when the diff is published until the recommendations are displayed. We monitor the 90th and 99th percentile as our goal is to reduce the time it takes for an author to see candidate reviewers
Clicks
The number of times that an author selects one of the recommended reviewers.
Workload
We measure the amount of time a candidate reviewer has spent looking at diffs in the last week.
@@ 4. Recommender Design and Evaluation Metrics
Table 2. The Features for RevRecV2 sorted by the feature importance in the model deployed in Section 5. We only include features that can be calculated with low latency, we use the feature categories suggested by Al-Zubaidi et al. (Al-Zubaidi et al., 2020).
Category Description Feature Importance
Familiarity with Author
@@
The time the candidate spent in the diff review tool looking at the author’s diffs in the past 90 days
33.89%
Review Participation Rate
Number of diffs the candidate was assigned as reviewer. At Meta reviewers are assigned to many diff either directly or as part of a team or review group.
31.28%
Reviewing Experience
Number of diffs the candidate rejected
11.56%
Reviewing Experience
Number of diffs the candidate commented or acted on as a reviewer
10.10%
Familiarity with Author
@@
The time the candidate spent in the diff review tool looking at the author’s diffs when they were explicitly assigned as a reviewer in the past 90 days
6.69%
Code Ownership
Number of diffs the candidate authored in the past which touched any of the files in the diff.
4.91%
Reviewing Experience
Number of diffs the candidate commented on
1.57%
Review Participation Rate
@@
Number of diffs the candidate resigned from. When a reviewer does not feel qualified to review a diff they can resign from that review.
0.00%
Having reviewed the literature on code review recommendation, we created a recommender that would suggest the reviewer that was most likely to perform a review, i.e. make a comment or do an action such as accepting a diff. Table 2 describes each of the features that we included across the categories of metrics found in the literature: code ownership, reviewing experience, reviewer participation, and familiarity with the author. We first develop a recommender considers all the prior features but excludes workload because the results of factoring workload are mixed, i.e. it worked at Microsoft but did not work at Ericsson. Instead, we weight the top ranked reviewers by their workload to understand if the approach that worked at Microsoft works at Meta.
4.1. Design for RevRecV2

### v2

@@ Abstract.
Aim. The code review team at Meta is continuously improving the code review process. In this work, we report on three randomized controlled experimental trials to improve code reviewer recommendation.
Method. To evaluate the recommenders, we conduct three A/B tests which are a type of randomized controlled experimental trial. The unit is either the code diff (Meta’s term for a pull-request) or all the diffs that an author creates during the experimental period. We set goal metrics, i.e. those we expect to improve, and guardrail metrics, those that we do not want to negatively impact, i.e. analogous to safety metrics in medical trials. We test the outcomes using a t-test, Wilcoxon test, or Fisher test depending on the type of data.
Expt 1. We developed a new recommender, RevRecV2, based on features that had been successfully used in the literature and that could be calculated with low latency. In an A/B test on 82k diffs in Spring of 2022, we found that the new recommender was more accurate and had lower latency. The new recommender did not impact the amount of time a diff was under review. The results allowed us to roll-out the recommender in Summer of 2022 to all of Meta.
Expt 2. Reviewer workload is not evenly distributed, our goal was to reduce the workload of top reviewers. Based on the literature, and using historical data, we conducted backtests to determine the best measure of reviewer workload. We then ran an A/B test on 28k diff authors in Winter 2023 on a workload balanced recommender, RevRecWL. Our A/B test led to mixed results. When a low workload reviewer had reasonable expertise, authors selected them, however, the top recommended low workload reviewer was often not selected. There was no impact on our guardrail metrics of the amount of time to perform a review. This workload balancing replaced the recommender from the first experiment as the recommender in production at Meta.
Expt 3. Engineers at Meta often select a team rather than an individual reviewer to review a diff. We suspected the bystander effect might be slowing down reviews of these diffs because no single individual was assigned the review. On diffs that only had a team assigned, we randomly selected one of the top three recommended reviewers to review the diff with BystanderRecRnd. We conducted an A/B test on 12.5k authors in Spring 2023 and found a large decrease in the amount of time it took for diffs to be reviewed. We did not find that reviewers rushed reviews. The results were strong enough to roll this recommender out to all diffs that only have a team assigned for review.
Implications. Aside from the direct findings from our work, our findings suggest there can be a discrepancy between historical back-testing and A/B test experimental findings, and that more A/B tests are necessary to test recommenders in production. Outcome measures beyond accuracy are important. This is especially true in understanding how recommenders change a reviewer’s workload. We also see that the latency in displaying a recommendation can have a large impact on how often authors select recommendations making the reporting of latency an important metric for future work.
@@ 1. Introduction
Code review is a critical activity in the software development lifecycle that ensures that at least one other experienced developer examines the code for defects and other issues. Code review has evolved from the inspection of massive completed artifacts (Fagan, 1976) to the continuous code review of the difference between the changed files in the form of pull-requests, patchsets, and diffs (Rigby and Bird, 2013; Bacchelli and Bird, 2013). A key part of the success of code review is the reviewer who both learns from the change and also suggests improvements. Finding an optimal code reviewer in OSS and a large organization is a complex task, hence substantial research has been devoted to code review recommenders which are tools that suggest who should review the code. At Meta the reviewers or review teams are recommended to the author who selects the best person or team. The author can also assign reviewers who are outside of the recommendations. Much of that literature uses the assumption that the actual code reviewer was an optimal choice to evaluate the accuracy of the recommenders on historic data. Instead, we run A/B tests of newly developed recommenders to assess if the modifications made to recommenders have the desired effects, goals metrics, and no undesired side-effects, guardrails metrics.
@@
At Meta, we are constantly evolving our code review system. In this work, we describe a series of experimentally validated improvements that we made to our code reviewer recommenders. Our starting point was developed in 2018 and we refer to it as RevRecV1. To address latency and accuracy issues with RevRecV1, we developed RevRecV2 that is based on a literature review of successfully used features. Building on RevRecV2, we use measures of reviewer workload to spread review work and develop RevRecWL. Finally, we combat the bystander effect when a group is assigned to do a review with BystanderRecRnd. One of our major contributions is to experiment with our recommenders in production and to conduct randomized controlled experimental trials (A/B test) on each recommender across hundreds of thousands of reviews and 10’s of thousands of engineers. We provide an overview of the motivation for each recommender and briefly summarize the results.
@@
Expt 1. Accuracy and latency improvements with RevRecV2. On the code review team’s feedback group, some engineers complained that they were being incorrectly recommended for reviews. When we investigated the current recommender, RevRecV1, we found it was purely based on who had last modified the lines in the current diff (i.e. blame lines), and it would suggest reviewers who had moved onto other projects because they had edited the files in the past. We reviewed the literature and developed a recommender that took the most promising features into account.
@@
Result summary: In an A/B test, our recommended reviewers were more likely to be the ones actually conducting the review with a 14.19 percentage point improvement over the previous model. The recommendations were also displayed more quickly with a 14 times reduction in latency at the 90th percentile. The recommendations were selected by authors 27% more often with the new model. We did not see any statistically significant regressions in our guardrail metrics including the review cycle time and the reviewer viewing time. In Summer 2022, RevRecV2 was rolled out to 100% of engineers.
@@
Expt 2. Reviewer workload balancing with RevRecWL. The number of reviews that each engineer does is known to be skewed (Rigby et al., 2014; Hajari, 2022; Al-Zubaidi et al., 2020). At Microsoft (Asthana et al., 2019) and Ericsson (Strand et al., 2020), they implemented a recommender that weighted the expertise based recommender score by the workload of the candidate reviewer. We replicate these works and experiment with novel workload measures, including the number of scheduled meetings to understand if workload balancing is effective.
@@
Result summary: When a reasonable candidate with lower workload was available, we saw a large reduction in workload. When an inappropriate candidate was recommended, simply because they had low workload, the author looked down the list for a more experienced developer, and we saw a drop in the top ranked candidate being selected. We did not see any statistically significant change in our guardrails: review cycle time or latency. In Winter of 2023, RevRecWL was rolled out to 100% of engineers.
@@
Expt 3. Reducing the bystander effect with BystanderRecRnd. The bystander effect occurs when nobody is explicitly assigned responsibility for a situation and no individual takes action or they are delayed in taking action (Fischer et al., 2011). Beyond noting that it may be an issue in code review (Meneely et al., 2014), we did not find any research in the software engineering literature. At Meta teams and groups of developers can be assigned a review without any explicit individual. To reduce the bystander effect, we use our recommenders to explicitly assign a reviewer.
@@
Result summary: At Meta, we randomly assigned one of the top three recommended reviewers to review the diff and found a drop of -11.6% in TimeInReview with no regressions in our guardrail metrics. In Spring of 2023, we rolled BystanderRecRnd out to 100% of diffs that only had a team assigned to do the review.
The remainder of this paper is structured as follows. In Section 2, we describe Meta and its code review process. In Section 3, we review the literature to identify the features that have been successfully used to recommend reviewers, and the design of our recommenders. In Section 4, we describe our experimentation methodology. In Section 5, we update the recommender at Meta to include these features and experiment with a new model in production and report results for Expt 1. In Section 6, we weight the recommender by the workload of the candidate reviewers in an attempt to balance workload and report results for Expt 2. In Section 7, for Expt 3, we reduce the impact of bystander effect when no individual is assigned a review by randomly assigning one of the recommended reviewers. In the final sections of the paper, we discuss threats to validity in Section 8, position our work in the literature in Section 9, and conclude the paper in Section 10.
@@ 2. Background and Code Review Process
Meta is a large online services company that works across the spectrum in the communications, retail, and entertainment industries. Meta has tens of thousands of employees with offices spread across the globe (North America, Europe, Middle East and Asia). Meta has its own dedicated infrastructure teams where the tools used are a mix of commercial and in-house developed systems. We study code reviews from the projects using Phabricator  [^1] Phabricator is an open source project available at https://www.phacility.com/phabricator/ which includes user facing social network products and virtual and augmented reality projects as well as software engineering infrastructure, such as calendar, tasks, and release engineering tooling.
@@ 2.1. Code review process
Code review at Meta follows a similar structure to the contemporary practices used at Google (Sadowski et al., 2018) and Microsoft (Bacchelli and Bird, 2013), and we briefly summarize the practice here. Figure 1 provides an example of a diff review in Phabricator with portions redacted and using mock names. The author fills in a number of fields including the title and a detailed summary of the diff change. A test plan is provided that details which steps the author took to validate their changes, and provides reviewers a guide to test the change themselves. Line-by-line highlighting is used to show additions and removals, and inline comments are used to indicate any compilation or linting issues. The diff will also show the output of tests that were required to run based on the files included in the change.
To start a code review, the authors must select one or more reviewers and/or reviewer groups. The author can explicitly type in names, or select recommended reviewers. The latency is a key consideration for a recommender because slow recommendations will not appear in time to be selected by the author. Once selected, the reviewer or a group of reviewers are notified via email or chat. The reviewers provide feedback, ask questions, and make suggestions, which should be addressed by the author. There may be multiple rounds of feedback and changes before a reviewer accepts/rejects the diff. Meta uses a single track/monorepo development model, so once the diff is accepted by at least one reviewer it will be integrated for further testing and ultimately “shipped“ into production.
@@ 2.2. Reviewer Recommenders at Meta
The reviewer recommender used in production at Meta before we began our work was developed in 2018. For this study it will be our starting comparison point and we will refer to it as RevRecV1. RevRecV1 considered the following attributes: the blame lines i.e. the person who has authored the most lines in a file is an expert, authorship frequency i.e. the person who has made the most changes to a file is an expert, and review frequency i.e. the number of times a reviewer has accepted the changes to a file. RevRecV1 is trained using a genetic algorithm, an ML approach popular in 2018. While the recommender has been effective and makes thousands of recommendations every day, the code review team identified a number of issues with it. First, many developers complain about incorrect recommendations on the internal feedback group. Second, blame is a computationally expensive command that can delay the loading of the code review page for a diff. For these reasons, the code review team decided to test alternative strategies for recommending reviewers to improve accuracy and reduce latency.
3. Literature, Features, and Recommender Design
We reviewed the literature to identify the reviewer recommendation approaches and model features that were most successful in existing recommenders. Like Microsoft (Asthana et al., 2019; Zanjani et al., 2016) and Ericsson (Strand et al., 2020), we cannot use features in production that would add latency to our recommender. We also found that very few recommenders that are studied in the literature report latency and even fewer are actually deployed into production. In this section, we update Al-Zubaidi et al.’s (Al-Zubaidi et al., 2020) literature review and use their categories: code ownership, reviewing experience, review participation rate, and familiarity with the author. Table 1 shows the features we selected for RevRecV2. In Section 3.2, we discuss how we weight RevRecV2 by different types of workload to create RevRecWL.
@@ 3. Literature, Features, and Recommender Design
The influence of code ownership on code quality has been extensively investigated in the literature(Bird et al., 2011; Rahman and Devanbu, 2011; Foucault et al., 2015; Thongtanunam et al., 2016) with recommenders to identify expert developers for a particular part of code or task (Mockus and Herbsleb, 2002). Researchers have used a wide range of granularity, from lines (Fritz et al., 2007; Girba et al., 2005; Rahman and Devanbu, 2011) to modules (Bird et al., 2011), to estimate ownership of developers. Studies on code review find that code owners are usually selected to review changes (Sadowski et al., 2018; Bacchelli and Bird, 2013; Greiler et al., 2016). In this work, we quantify ownership as the number of files under review that a candidate reviewer has modified in past diffs.
@@
Thongtanunam et al. (Thongtanunam et al., 2016) recommend reviewers based on the files that an engineer has reviewed. Intuitively, engineers who have reviewed the files in the diff in the past, will be good candidate reviewers. Zanjani et al. (Zanjani et al., 2016) builds upon this notion of reviewer expertise by adding the number of comments a candidate reviewer has made on the files under review. In this work, we count the number of comments as well as other reviewer actions such as accepting or rejecting the diff.
@@
Moving beyond code ownership, reviewer recommenders have begun to consider relationships among the reviewer and the author (Bosu et al., 2016). The simplest metric is the number of times a reviewer has commented on an author’s diffs in the past. More sophisticated usages include social relationship graphs (Yu et al., 2016). With 10’s of thousands of developers, and thousands of diffs every day, social network graphs are not performant. In this work, we suggest a novel usage of how long a reviewer looks at an author’s diffs: the time a reviewer spent looking at the diffs the author wrote in the code review tool. While Meta has sophisticated tracking of how long an engineer is looking at a webpage and uses this to capture how long a reviewer looks at a review, other projects can easily add this to the review tool. For example, the SmartBear code review tool has been tracking this information since 2006 (Cohen, 2006) to identify reviewers that looked at the change for too short a period of time given the size of the diff, i.e. “rubberstamping”. The web abounds with code to monitor time spent on a webpage. [^2] A JavaScript library that tracks time spent on a page https://stackoverflow.com/questions/4667068/how-to-measure-a-time-spent-on-a-page
@@
Recent works have incorporated the reviewer workload into reviewer recommendation. Early works quantified the relationship between time waiting for a review and review workload. Baysal et al. (Baysal et al., 2013) studied WebKit to understand the impact of metrics on review time. They found that when reviewers had longer queues of waiting reviews, reviews tended to take longer. They also found that more active reviewers had quicker response times than less active reviewers. Studying four open source projects, Ruangwan et al. (Ruangwan et al., 2019) found prior participation and prior experience were the best predictor of whether a reviewer will respond to a review request. The familiarity of the author and reviewer had a minor impact on participation.
@@
Microsoft weighted their code reviewer recommender, cHRev, by the number of open reviews a candidate, WhoDo, had and found between a 14% and 21% improvement in the time taken for reviews depending on the project (Asthana et al., 2019). In contrast, Strand  et al. (Strand et al., 2020) evaluates a workload aware recommender at Ericsson. In a backtest they found that a recommender aware of the file change history and the level of recent workload activity was able to make recommendations with an MRR of 0.37 or an average rank of 2.7. Despite these reasonable recommendations, after manually evaluating the performance of the recommender on 47 changes they determined that it did not reduce lead time or workload for code review and they decided to not roll-out the recommender.
@@
Table 1. The Features for RevRecV2 and their importance in the model deployed in Section 5. We only include features that can be calculated with low latency, we use the feature categories suggested by Al-Zubaidi et al. (Al-Zubaidi et al., 2020). Workload is only part of RevRecWL and is not included in this table.
Category Feature Description Importance
@@
Familiarity with Author time spent The time the candidate spent in the diff review tool looking at the author’s diffs in the past 90 days 33.89%
@@
Familiarity with Author explicit time spent The time the candidate spent in the diff review tool looking at the author’s diffs when they were explicitly assigned as a reviewer in the past 90 days 6.69%
Review Participation Rate assigned reviews Number of diffs the candidate was assigned as reviewer. At Meta reviewers are assigned to many diff either directly or as part of a team or review group. 31.28%
@@
Review Participation Rate resigned diffs Number of diffs the candidate resigned from. When a reviewer does not feel qualified to review a diff they can resign from that review. 0.00%
Reviewing Experience rejected diffs Number of diffs the candidate rejected 11.56%
Reviewing Experience reviewer comments Number of diffs the candidate commented or acted on as a reviewer 10.10%
Reviewing Experience total comments Number of diffs the candidate commented on 1.57%
Code Ownership authored diffs Number of diffs the candidate authored in the past which touched any of the files in the diff. 4.91%
3.1. Design and Feature Importance for RevRecV2